OpenStudy (anonymous):

How to solve this Rational Algebraic Expression 5b/3b^2+2b

OpenStudy (saifoo.khan):

try factoring the denominator.

OpenStudy (anonymous):

and then?

OpenStudy (saifoo.khan):

\[\frac{5b}{b(3b+2)}\]\[\frac{5 \cancel{b}}{\cancel{b}(3b+2)} \]

OpenStudy (saifoo.khan):

Finally, \[\frac{5}{3b+2}\]

OpenStudy (anonymous):

Nice.!